Please I hope some one can help me, I am not doing to well in my algebra class.I have three questions I am having trouble with.I want to thank you for all your help in advance.
1. Solve the system of equations using the substitution method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
-3x + y = 1
5x + 2y = -4
2. Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
-7x + y = 8
2x – y = 2
3. Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
3x – 2y = -7
-9x + 6y = 21

Answer by stanbon(75887)

1. Solve the system of equations using the substitution method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
-3x + y = 1
5x + 2y = -4
----------------
y = 3x+1
Substitute into the 2nd equation
5x+2(3x+1) = -4
11x+2 = -4
x = -6/11
----------------
Substitute that into y = 3x+1:
x = 3(-6/11) + 1
x = -18/11 + 11/11
x = -7/11
----------------
2. Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
-7x + y = 8
2x – y = 2
-----
Add the equation to get:
-5y = 10
y = -2
--------
Substitute into 2x-y = 2 to solve for "x":
2x-2 = -2
x = 0
---------------------------
3. Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
3x – 2y = -7
-9x + 6y = 21
-----
Multiply thru the 1st equation by 3:
9x -6y = -21
Add to the 2nd equation:
0= 0
-----------
This means the two equations are dependent; one is a multiply of the other.
The solution is all points where 3x-2y=-7
===========================================
Cheers,
Stan H.